# Synthia: Novel Concept Design with Affordance Composition

### v2

@@ B.2 Automatic Evaluation
Appendix C Human Evaluation Details
@@
Human Evaluation Instruction Objective The goal of this evaluation is to assess the quality of novel concepts that integrate multiple affordances into a single, coherent design. Affordance refers to the functional properties of an object or its components. For example, a sofa affords the function of sitting, while its legs provide the function of support. As an annotator, you will evaluate the given concepts based on four key metrics: faithfulness, novelty, practicality, and coherence. Each metric is defined below, along with its respective scoring criteria. Evaluation Criteria Faithfulness (Does the concept effectively integrate the specified affordances?) This metric assesses whether the generated concept successfully incorporates all provided affordances in a meaningful and functional manner. Scoring Scale: 5 – Fully integrates all specified affordances, demonstrating a well-designed and fully functional object. 4 – Incorporates all affordances with minor inconsistencies or slight missing details. 3 – Partially fulfills the affordances; some functionalities are present but not well-integrated or consistent. 2 – Struggles to meet the provided affordances; key functionalities are missing or poorly combined. 1 – Does not incorporate the specified affordances; functionalities are entirely missing, irrelevant, or nonsensical. Novelty (To what extent does the concept demonstrate originality and innovation?) This metric evaluates the uniqueness and creative appeal of the design, considering whether it introduces novel elements that would intrigue or surprise users. Scoring Scale: 5 – Highly innovative and unique; presents a strikingly original concept that is engaging and thought-provoking. 4 – Contains clear novel elements, demonstrating creative thought and originality. 3 – Moderately novel; some unique aspects are present, but the overall concept remains relatively conventional. 2 – Limited novelty; the design appears simplistic, derivative, or lacking in creativity. 1 – Entirely unoriginal and uninspiring, closely resembling existing designs with no innovative aspects. Practicality (Is the design feasible and suitable for real-world use?) This metric assesses whether the concept is functionally viable and aligned with human preferences and usability considerations. Scoring Scale: 5 – Highly practical and user-centered; seamlessly functional and feasible for real-world applications. 4 – Mostly practical; minor limitations exist but do not significantly hinder usability. 3 – Somewhat practical; contains notable flaws or unrealistic elements that may limit real-world applicability. 2 – Largely impractical; significant design flaws make real-world usability unlikely. 1 – Entirely impractical and non-functional; does not align with human preferences or feasibility constraints.
Human Evaluation Instruction (continued) Coherence (Does the image clearly depict a single, distinct object?) This metric evaluates whether the design presents a singular, well-defined object, free from extraneous elements that may obscure its intended functionality. Scoring Scale: 5 – The image clearly and exclusively depicts a single object that integrates all specified affordances without any distractions. 4 – The primary object is distinct and well-defined, though minor background elements may be present without significantly detracting from clarity. 3 – The main object is identifiable, but additional elements in the image introduce some visual or conceptual distractions. 2 – The image contains multiple objects, making it difficult to distinguish the intended primary object or missing at least one affordance. 1 – The image primarily features multiple objects, with affordances spread across different elements rather than a unified concept, making it unclear what the primary object is. Final Instructions Please evaluate each concept independently based on the above criteria. Assign a score for each metric according to the provided descriptions. If a concept does not fit neatly into the scoring categories, use your best judgment to determine the most appropriate score. Your evaluations will contribute to assessing the effectiveness of novel concept generation and help improve future designs. Thank you for your participation.
Appendix D Experiments
D.1 Dataset
  • Existing Concept Images: For each existing concept in our ontology, we collect a dataset of 60 images from external platforms including Google Images and iStock. To ensure that the dataset is object-centric and minimizes noise, we filter out low-quality images using CLIP model (Radford et al., 2021). Specifically, we compute the similarity between the image embeddings and the text embeddings of the "a photo of {concept name}", selecting top-5 images with the highest similarity scores for each concept used as negative constraints.
  • Generated Novel Concept Images: With our affordance sampling, we uniformly sample 600 affordance pairs among 82K possible pairs for fine-tuning. For the test dataset, we select 500 affordance pairs among the ones not used for fine-tuning. We use the generated images from the sampled affordances for fine-tuning and evaluation. The overall statistics can be found in Table  2.
Dataset Fine-tuning Inference
Existing Generated Generated
# Concepts 772 600 500
# Images 3860 1800 500
Table 2: Statistics of the datasets.
D.2 Baseline Methods
  • Stable Diffusion (Esser et al., 2024) is a strong baseline model for high-fidelity image synthesis, which is built on a diffusion-based framework. In this work, we leverage the pre-trained stable-diffusion-3.5-large model as the foundation model for the text-to-image task to generate a novel concept. Due to the limited context window length, we input only the positive constraints that contain the desired affordances, omitting the negative constraints associated with existing concepts.
  • Kandinsky (Arkhipkin et al., 2023; Vladimir et al., 2024) model serves as another strong baseline model for comparison. We utilize the pre-trained Kandinsky 3.0 model without any finetuning as the baseline, aligning it with the foundation model used in our proposed framework. This approach ensures a consistent starting point for evaluation and a fair comparison. This baseline allows us to effectively demonstrate the impact of our proposed training framework by comparing performance before and after the fine-tuning process.
  • ConceptLab (Richardson et al., 2024) is a state-of-the-art framework designed for creative concept generation, which leverages an innovative approach that formulates the generation problem as an optimization process over the output space of the diffusion prior. It adopts a similar input format to the one used in our setting. We follow the finetuning process in the original framework, applying our training data to generate novel concepts. We then compare the generation quality during the inference time.
Figure 6: Results of the absolute automatic evaluation across different number of training data in Synthia.
@@ D.2 Baseline Methods
Figure 7: Results of the relative automatic evaluation. We compare the quality of concepts generated from our models with ones from existing T2I models. Numbers indicate the percentage (%) of baseline model wins, ties, and DALLE model wins.
D.3 Hyper-parameter Settings
To fine-tune the diffusion model with the constructed dataset from DALL-E, we use the optimizer AdamW (Loshchilov and Hutter, 2019) with a learning rate of $\eta=\{5*10^{-6},10^{-6},5*10^{-7}\}$ to avoid the catastrophic forgetting problem. During the curriculum learning, we split the dataset into three groups based on their difficulty and trained 20 epochs for each group. For the weight factor in the triplet loss, we set $\gamma=\{0,0.2,0.5,0.8,1\}$. We finetune the UNet part in the pre-trained model and freeze the weights of other components in the diffusion model.
Appendix E Additional Experimental Results
E.1 Effectiveness of affordance-based textual prompt
@@
Table 3: Results of the automatic evaluation with three positive affordances. We compare our method with baseline models. For each metric, a higher number indicates a better performance, where the absolute score ranges between 0 and 5.
@@
Table 4: Results of the automatic evaluation with four positive affordances. We compare our method with baseline models. For each metric, a higher number indicates a better performance, where the absolute score ranges between 0 and 5.
@@
Table 5: Results of the automatic novelty evaluation with different distances. We compare our method with baseline models. For each metric, a higher number indicates a better performance, where the absolute score ranges between 0 and 5.
@@
Table 6: Results of the automatic evaluation with and without curriculum learning. For each metric, a higher number indicates a better performance, where the absolute score ranges between 0 and 5.
Figure 8: Results of the human evaluation.

### v4

@@ B.5.2 Relative Automatic Evaluation
B.6 Human Evaluation
@@
Human Evaluation Instruction Objective The goal of this evaluation is to assess the quality of novel concepts that integrate multiple affordances into a single, coherent design. Affordance refers to the functional properties of an object or its components. For example, a sofa affords the function of sitting, while its legs provide the function of support. As an annotator, you will evaluate the given concepts based on four key metrics: faithfulness, novelty, practicality, and coherence. Each metric is defined below, along with its respective scoring criteria. Evaluation Criteria Faithfulness (Does the concept effectively integrate the specified affordances?) This metric assesses whether the generated concept successfully incorporates all provided affordances in a meaningful and functional manner. Scoring Scale: 5 – Fully integrates all specified affordances, demonstrating a well-designed and fully functional object. 4 – Incorporates all affordances with minor inconsistencies or slight missing details. 3 – Partially fulfills the affordances; some functionalities are present but not well-integrated or consistent. 2 – Struggles to meet the provided affordances; key functionalities are missing or poorly combined. 1 – Does not incorporate the specified affordances; functionalities are entirely missing, irrelevant, or nonsensical. Novelty (To what extent does the concept demonstrate originality and innovation?) This metric evaluates the uniqueness and creative appeal of the design, considering whether it introduces novel elements that would intrigue or surprise users. Scoring Scale: 5 – Highly innovative and unique; presents a strikingly original concept that is engaging and thought-provoking. 4 – Contains clear novel elements, demonstrating creative thought and originality. 3 – Moderately novel; some unique aspects are present, but the overall concept remains relatively conventional. 2 – Limited novelty; the design appears simplistic, derivative, or lacking in creativity. 1 – Entirely unoriginal and uninspiring, closely resembling existing designs with no innovative aspects. Practicality (Is the design feasible and suitable for real-world use?) This metric assesses whether the concept is functionally viable and aligned with human preferences and usability considerations. Scoring Scale: 5 – Highly practical and user-centered; seamlessly functional and feasible for real-world applications. 4 – Mostly practical; minor limitations exist but do not significantly hinder usability. 3 – Somewhat practical; contains notable flaws or unrealistic elements that may limit real-world applicability. 2 – Largely impractical; significant design flaws make real-world usability unlikely. 1 – Entirely impractical and non-functional; does not align with human preferences or feasibility constraints. Coherence (Does the image clearly depict a single, distinct object?) This metric evaluates whether the design presents a singular, well-defined object, free from extraneous elements that may obscure its intended functionality. Scoring Scale: 5 – The image clearly and exclusively depicts a single object that integrates all specified affordances without any distractions. 4 – The primary object is distinct and well-defined, though minor background elements may be present without significantly detracting from clarity. 3 – The main object is identifiable, but additional elements in the image introduce some visual or conceptual distractions. 2 – The image contains multiple objects, making it difficult to distinguish the intended primary object or missing at least one affordance. 1 – The image primarily features multiple objects, with affordances spread across different elements rather than a unified concept, making it unclear what the primary object is. Final Instructions Please evaluate each concept independently based on the above criteria. Assign a score for each metric according to the provided descriptions. If a concept does not fit neatly into the scoring categories, use your best judgment to determine the most appropriate score. Your evaluations will contribute to assessing the effectiveness of novel concept generation and help improve future designs. Thank you for your participation.
Appendix C Additional Experimental Results
C.1 Results of Relative Automatic Evaluation
As described in Sec 4.3, we perform relative evaluation following the protocol B.5.2, comparing each method s generated concepts against those from DALL-E. We show that Synthia achieves higher win rates than the baselines in all metrics when evaluated against DALL-E. Moreover, we conduct direct pairwise comparisons between Synthia and the baselines. As shown in Fig. 7, Synthia outperforms all baselines across every metric, demonstrating its superior ability to design novel concepts that are both visually novel and functionally coherent, given a set of desired affordances.
@@ C.1 Results of Relative Automatic Evaluation
Figure 7: Results of the relative automatic evaluation. We compare the quality of concepts generated from our models with ones from existing T2I models. Numbers indicate the percentage (%) of baseline model wins, ties, and Synthia model wins.
C.2 Results of Human Evaluation
In addition to the absolute inter-annotator agreement (IAA) reported in Sec 4.3, we further visualize the agreement between individual human annotators. As shown in Fig. 8, the overall agreement is high, supporting the reliability and validity of our human evaluation results.
Figure 8: Results of the human evaluation.
C.3 Effectiveness of Affordance-based Textual Prompt
@@
Table 4: Results of the automatic evaluation with three positive affordances. We compare our method with baselines. For each metric, a higher number indicates a better performance, where the score ranges between 0 and 5.
@@
Table 5: Results of the automatic evaluation with four positive affordances. We compare our method with baselines. For each metric, a higher number indicates a better performance, where the score ranges between 0 and 5.
C.4 Effectiveness of Affordance Sampling
@@
Table 6: Results of the automatic novelty evaluation with different distances. We compare our method with baselines. For each metric, a higher number indicates a better performance, where the score ranges between 0 and 5.
C.5 Effectiveness of Curriculum Learning
@@
Table 7: Results of the automatic evaluation with and without curriculum learning. For each metric, a higher number indicates a better performance, where the score ranges between 0 and 5.
C.6 More Visual Examples
We include additional visual examples comparing the baseline methods with Synthia (Ours). As illustrated in Figure 9, our model outperforms in generating concepts that align closely with the input affordance pairs. Moreover, Synthia consistently produces outputs with higher quality across 4 key dimensions, including faithfulness, novelty, practicality, and coherence.
Figure 9: More generated examples from the baseline methods and Synthia (ours).
Appendix D Relevance to NLP domain
While our work focuses on Text-to-Image (T2I) generation, it is fundamentally grounded in Natural Language Processing (NLP) domain. Specifically, we leverage NLP methods to extract concept-affordance associations from commonsense knowledge bases and incorporate the structured knowledge into the image generation process. Our interdisciplinary approach not only expands the horizons of text-to-image generation, but also advances core NLP research by demonstrating how linguistic and semantic understanding can inform and enrich multimodal systems. This contribution aligns closely with the NLP community growing interest in multimodal AI and the integration of language with other modalities.
Ethical Consideration